OpenStudy (anonymous):

How do you find the point on a line and the slope from an equation in standard form: Ax+By=C?

Nnesha (nnesha):

to find the slope you should solve for y first

Nnesha (nnesha):

slope intercept form equation is should look like this y=mx+b here m is slope and y is y-inrercept

OpenStudy (anonymous):

Well the slope is equal to -A/B

OpenStudy (anonymous):

for y=mx+b the slope is m right?

OpenStudy (anonymous):

Y-intercept is C/B

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Yes, the slope (m)= -A/B

Nnesha (nnesha):

yes right which always with x

Nnesha (nnesha):

to solve this equation Ax+By =C first subtract ax both side

Nnesha (nnesha):

then divide both side by B

OpenStudy (anonymous):

okay then divide b to get y by itself

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

\[\Large Ax+By=C\] \[\Large By= -Ax+C\] \[\Large y= \frac{ -A }{ B }x+\frac{ C }{B}\]

Nnesha (nnesha):

let me know what you get :)

Nnesha (nnesha):

well nvm you already got answer

OpenStudy (anonymous):

where \[\Large y=mx+b\] is the same as \[\Large y= \frac{ -A }{ B }x+\frac{ C }{B}\]

OpenStudy (anonymous):

Ohkay this makes sense

OpenStudy (anonymous):

\[\Large m=\frac{ -A }{ B }\] \[\Large b=\frac{ C }{ B}\]